A laser beam is incident on 2 slits with a separation of 0.200nm and a screen is placed 5.00m from the slits. An interference pattern appears on the screen. If the angle from the center fringe to the first bright fringe to the side is 0.181 degree, what is the wavelength of the laser light?


1
Expert's answer
2021-02-22T10:28:23-0500

Given parameters are-

slit width, "d=0.2mm=0.2\times 10^{-3}m"

Angle, "\theta=0.181^{\circ}"

order, "n=1"


Using the formula,

"n\lambda=dsin\theta"

"\Rightarrow 1\times \lambda=0.2\times 10^{-3} sin 0.181^{\circ}\\\Rightarrow \lambda=6.318\times 10^{-7}m"


Hence The wavelength is "6.318\times 10^{-7}m"